OpenStudy (anonymous):

solve the equations for the variable log2 x + log2 (x-2)-3=0

OpenStudy (anonymous):

x = 4 or -2

OpenStudy (lalaly):

\[\log_2 x(x-2) =3\]\[2^3=x(x-2)\]\[x^2-2x-8=0\]\[(x-4)(x+2) =0\]

OpenStudy (lalaly):

x=4 or -2